Grade 8General Physics

a ball of mass m kg is thrown vertically upwards. Another ball of mass 2m is thrown at angle θ with the vertical. Both stays in air for the same period . the heights reached by them are in the ratio of
(a) 2:1
(b) 1:!
© 1:cosθ
(d) 1:sinθ

hari
8 Years agoGrade 8

1 Answer

Arun
8 Years ago
height obtained by an object thrown up with u velocity (h) = u^2/2g
 
Hence, ratio of the heights (h1/h2) = u1^2/u2^2
Also, since they remain in the air for same duration. 
So, using first law of motion v = u + at
v = 0, t = -u/a, Now here, since the a is same as equal to g and t is same, so u needs to be same.